150 R I you 70 R qov go n ㅈ 피need a break down of how to find magnitude and direction using kirschoffs law, thanks!
Answer #1

2 Lov Iz A 702 qov qon Isor I, + I ₂ = I3 -(1) (At point A) using KVL in toplowy TIIS 150 I2 + 70 Iz = 40 -621 using KU L in